Suppose that you want to purchase a used car for $12000. You currently have $8000. If you invest your $8000 at an annual interest rate of 3%, compounded quarterly, how long will it take (in years) before you can purchase the car?
Answer by Theo(13342)

the price of the car is 12,000.
you have 8,000.
the annual interest rate is 3% compounded quarterly.
how many years before you can purchase the car?

the formula to use is f = p * (1 + r) ^ n

f is the future value.
p is the present value.
r is the interest rate per time period.
n is the number of time periods.

since you are compounding the interest rate quarterly, your interest rate per quarter is equal to 3% / 4 which is equal to .75% expressed as a percent and .0075 expressed as a rate.
in the formula, you use the rate, not the percent.

n is also expressed in quarters of a year, so if you want to know the number of years, you would have to divide n by 4 as well, once you find it.

your formula becomes 12,000 = 8,000 * (1 + .0075) ^ n.

divide both sides of this equation by 8,000 to get 12,000 / 8,000 = (1 + .0075) ^ n.

take the log of both sies of this equation to get log(12,000/8,000) = log(1.0075^n).

since log(1.0075^n) is equal to n * log(1.0075), your formula becomes log(12,000/8,000) = n * log(1.0075).

divide both sides of this equation by log(1.0075) and you get log(12,000/8,000) / log(1.0075) = n.

solve for n to get n = 54.2644945.

replace n in the original equation with this to get 12,000 = 8,000 * (1.0075)^54.2644945.

the result is 12,000 = 12,000.

recall that n is in quarters, so divide 54.2644945 by 4 to get 13.56612362 years.

an alternate way of analyzing this is to use the formula f = p * (1 + r/c) ^ (n*c).

this formula assumes annual interest rate and number of years and incorporates the number of compounding periods per year into it.

using this formula, you would get 12,000 = 8,000 * (1 + .03/4) ^ (n*4).

you would ue the same procedure to get log(12,000 / 8,000) / log(1 + .03/4) = n*4.

you would solve for n * 4 to get n * 4 = 54.2644945.

you would then solve for n to get n = 13.56612362.

in the formula of f = p * (1 + r/c) ^ (n*c), .....

f = future value as before
p = present value as before
r = annual interest rate (not the annual interest rate percent).
n = number of years
c = number of compounding periods per year.

in the first formula of f = p * (1 + r) ^ n,
r is assumed to be per time period and n is assumed to be number of time periods.
the adjustment from years to time periods is done prior to using the formula.


in the second formula of f = p * (1 + r/c) ^ (n*c),
r is assumed to be per year and n is assumed to be number of years and c is the number of compounding periods per year.
the adjustment from years to time periods is incorporated as part of the formula.